OpenStudy (anonymous):

What is 243 written as a power of a number? A. 3^4 B. 5^3 C. 3^5 D. 81^3

OpenStudy (anonymous):

3^5

OpenStudy (whpalmer4):

\[3^4 = 3*3*3*3 = \]\[5^3= 5*5*5=\]\[3^5= 3*3*3*3*3=\]\[81^3=81*81*81=\]

rvc (rvc):

3^5

OpenStudy (anonymous):

thanks